Solve this system using elimination. If a single solution exists, write the solution as an ordered pair. Your answer will be an ordered pair, no solution, or infinitely many solutions.2x-y=0 and 3x+y=5

Answer by stanbon(75887)

Solve this system using elimination. If a single solution exists, write the solution as an ordered pair. Your answer will be an ordered pair, no solution, or infinitely many solutions.
2x-y=0
3x+y=5
----------
Add to get:
5x = 5
x = 1
---
solve for "y":
2x-y = 0
y = 2x
y = 2*1 = 2
----
Single solution: (1,2)
==============================
Cheers,
Stan H.
============